# Thought2Text: Text Generation from EEG Signal using Large Language Models (LLMs)

### v1

@@ 6 Results
Table 2 presents a comprehensive comparison of metrics across different models and setups. From these results, it is evident that the complete approach, denoted as ALL, consistently outperforms other setups across all evaluation metrics. As expected, chance-based baselines like and exhibit poor performance. For instance, the LLaMA3-8B_ALL model achieved a BLEU-N (N=1) score of 25.5%, Mistral-7B_ALL scored 26%, and QWEN2.5-7B_ALL reached 22.7%—all significantly higher than their respective chance scores under the chance setups. Notably, some models, particularly the LLaMa variants, show increased sensitivity to random input, leading to further reduction in the scores with chance setups.
@@ 6.1 Comparison with Stage 2 Omission (NO_STAGE2)
The ROUGE-N (N=1) score for LLaMA3-8B’s variant is 30.0%, whereas the variant achieves only 26.9%, indicating a significant improvement with the inclusion of Stage 2. The same trend is observed for other metrics like BLEU-N (N=1 and N=4), METEOR, and BERT Score, and is observed across other models as well. In terms of adequacy as measured by GPT-4, the variant stands out except for Qwen model which shows a tendency to copy the object and produce shorter sentences in the case of which is positively rated by GPT-4. Overall, the table demonstrates that incorporating Stage 2 (as detailed in Section 4) — which aligns EEG embeddings with image embeddings using a CLIP-based supervision strategy — contributes to higher-quality text generation. The consistently higher BERT scores further emphasize that Stage 2 helps maintains semantic consistency, which aligns with improved Fluency and Adequacy as assessed by GPT-4.
@@ 6.2 Generation performance without object labels in the input
While our initial hypothesis was that EEG embeddings—derived from noisy, multichannel EEG data—might not fully capture complex thoughts, thus requiring additional inputs like object labels, it is pleasantly surprising to observe that even without object labels, the models () perform comparably to the best models (). This underscores the effectiveness of aligning EEG embeddings with vision embeddings in Stage 1 and pretraining the LLM, particularly the projectors, in Stage 2 with vision embeddings.
@@ 6.3 Subject-wise Analysis
For this analysis, each subject’s EEG data is used to independently train and test the LLMs, simulating a personalized solution. The dataset comprises six subjects, allowing for individual analysis to evaluate the robustness of the approach across different participants.
As depicted in Figure 2, in the subject-wise analysis, the advantages of the complete approach () become even more prominent. Cross-subject and in-subject experiments consistently favor the configuration, with significant improvements in Adequacy score when Stage 2 is included, especially for LLaMA3-8B and Qwen2.5-7B models. These improvements, though numerically small, are crucial in the context of EEG data where every bit of alignment and finetuning matters due to its inherently noisy and sparse nature.
The configuration, which omits the essential alignment step between EEG embeddings and image embeddings, consistently demonstrates lower performance across subjects, as illustrated by the BLEU Unigram scores in Figure 2. This validates our hypothesis that direct fine-tuning without the warm-up provided by Stage 1 and Stage 2 is insufficient for EEG data. The challenge is further exacerbated by the inherent difficulties associated with EEG data, which, even when ethically collected and cleaned, still suffers from limited data availability and significant variability across sessions and subjects.
Subject-wise analysis is essential in practice due to the inherently private and sensitive nature of EEG data. Models like ours, designed for thought-to-text translation, must be developed and deployed within privacy-preserving settings. Creating a personalized solution without access to extensive cross-subject EEG data can be challenging.
To this end, our multi-stage approach demonstrates that pretraining on non-EEG data (such as images) and fine-tuning on small amounts of subject-specific EEG data opens new avenues for privacy-preserving personalized EEG-LLM model development. The efficacy of this approach, as opposed to direct fine-tuning following standard practices, shows that a combination of pretraining and fine-tuning can make personalized EEG models a reality, even with limited data.
@@ 6.4 Qualitative Inspection and Basic Error Analysis
For qualitative inspection, we compared the generated descriptions of images produced by both the GPT-4 and the Mistral-7B-v0.3 model ( and variants). Table 3 presents notable examples from our assessment of the model’s generated descriptions. In many instances, the approach with input of EEG + predicted object label generates highly accurate descriptions, as illustrated in positive examples 2 and 4 for the piano and pumpkin, respectively. However, in cases of misgeneration, we encounter not only significantly inaccurate outputs and hallucinations, as seen in examples 6, 7, and 8, but also instances of genuine confusion. For instance, the EEG signal for a flower is misidentified as a mushroom in example 5, leading to incorrect generation. This misidentification indicates potential areas for improvement in object classification.
However, a key advantage of the EEG-only approach is that, even when the predicted object label is incorrect, the model can still produce reasonably coherent descriptions, as seen in example 1. This highlights the robustness of the EEG embeddings in guiding the language model’s generation. Similar observations were made with other models based on LLaMa and QWEN architectures, further validating the consistency and reliability of our approach across various LLM frameworks. We acknowledge that the selected examples are purposefully chosen to illustrate different scenarios.
@@ 7 Conclusion and Future Work
In this paper, we introduced a novel approach to convert EEG signals into text, leveraging instruction-tuned LLMs fine-tuned with EEG data. Our method progresses through three stages: training an EEG encoder for feature extraction, fine-tuning LLMs on multimodal data, and further refining them with EEG embeddings for direct text generation from neural signals. Validation on a public EEG dataset demonstrates the efficacy of popular LLMs in translating visual thoughts into text. We observe significant performance improvements compared to chance evaluation, and our methodology, incorporating all stages, performed well in both cross-subject and in-subject analyses, as validated through quantitative evaluation. The qualitative evaluation further provided insights into various scenarios involving EEG signals and object labels versus EEG signals alone as inputs for generating text. These evaluations not only reinforce the effectiveness of our methodology for efficient text generation but also underscore the potential of utilizing EEG data alone to achieve the desired results. However, instances of misidentification that result in incorrect outputs reveal opportunities for further improvements in text generation.
Our future work will focus on optimizing model architectures, leveraging foundational pretrained EEG models on diverse dataset, improving EEG-text alignment through stage 2 training on large scale image datasets and diverse tasks such as optical character recognition, question answering and summarizing and exploring practical applications in healthcare and assistive technologies, marking a significant stride toward accessible "thoughts-to-text" systems.
@@
Extracting fine-grained information from EEG signals presents challenges due to high data-to-noise ratio and low spatial resolution. Despite these difficulties, EEG signals can still identify object categories, which can then be used with a generic prompt to aid in text generation. However, misclassification of similar-shaped objects (see the Appendix, Figure 3), such as mistaking a mushroom for a flower, underscores potential ambiguities in object recognition. Additionally, there were instances where the encoder generated unrelated descriptions, such as identifying a coffee maker as a computer or an elephant as a panda. Object classification accuracy varies among subjects (see the Appendix, Table 4) due to individual differences in interpreting images, leading to diverse EEG signal variations and increased prediction variance. One approach to address this is by training personalized models for each subject and assessing their performance. Additionally, implementing methods that enhance the generalizability of predictions across subjects could be explored. Another challenge is data scarcity; Deep Learning models typically require substantial data for training. Hence, acquiring more high-quality, multi-channel EEG data under controlled experimental conditions is crucial to reduce noise. Despite these challenges, our quantitative and qualitative findings demonstrate promising results. We believe that additional training on larger datasets and rigorous controlled experiments will significantly improve performance.
@@
For this work, we utilized anonymized open-source EEG data, acknowledging the sensitivity of EEG data and the imperative of ethical compliance. All experimental data used in our research were anonymized to protect participant privacy and uphold ethical standards.

### v2

@@ 6 Results
Table 2 presents a comprehensive comparison of metrics across different models and setups. From these results, it is evident that the complete approach, denoted as ALL, consistently outperforms other setups across all evaluation metrics. As expected, chance-based baselines like and exhibit poor performance. With our proposed methodology, the LLaMA-v3-8B_ALL model achieved a BLEU-N (N=1) score of 25.5%, Mistral-7B_ALL scored 26%, and Qwen2.5-7B_ALL reached 22.7%—-all significantly higher than their respective chance scores under the chance setups. In particular, some models, mainly the LLaMa variants, show increased sensitivity to random input, leading to further reduction in the scores with chance setups.
@@ 6.1 Comparison with Stage 2 Omission (NO_STAGE2)
The ROUGE-N (N=1) score for LLaMA3-8B’s variant is 30.0%, whereas the variant achieves only 26.9%, indicating a significant improvement when Stage 2 is added. The same trend is observed for other metrics like BLEU, METEOR, and BERT Score, and is observed across other models as well. In terms of adequacy as measured by GPT-4, the variant stands out except for Qwen model which shows a tendency to copy the object and produce shorter sentences in the case of which is positively rated by GPT-4. Overall, the table demonstrates that incorporating Stage 2 (as detailed in Section 4) – which aligns EEG embeddings with image embeddings using a CLIP-based supervision strategy – contributes to higher-quality text generation.
@@ 6.2 Generation performance without object labels in the input
While our initial hypothesis was that EEG embeddings – derived from noisy multichannel EEG data – might not fully capture complex thoughts, thus requiring additional inputs like object labels, it is pleasantly surprising to observe that even without object labels, the models () perform comparably with the best models (). This underscores the effectiveness of aligning EEG embeddings with vision embeddings in stage 1 and pretraining the LLM, particularly the projectors, in stage 2 with vision embeddings.
@@ 6.3 Subject-wise Analysis
For this analysis, each subject’s EEG data is used to independently train and test the LLMs, simulating a personalized solution. The dataset comprises six subjects, allowing for individual analysis to evaluate the robustness of the approach across different participants. As depicted in Figure 2, in the subject-wise analysis, the advantages of the complete approach (ALL) become even more prominent. Cross-subject and in-subject experiments consistently favor the ALL configuration, with significant improvements in Adequacy score when Stage 2 is included, especially for LLaMA3-8B and Qwen2.5-7B models. These improvements, though numerically small, are crucial in the context of EEG data where every bit of alignment and finetuning matters due to its inherently noisy and sparse nature.
The configuration, which omits the essential alignment step between EEG embeddings and image embeddings, consistently demonstrates lower performance across subjects, as illustrated by the BLEU Unigram scores in Figure 2. This validates our hypothesis that direct fine-tuning without the warm-up provided by Stage 1 and Stage 2 is insufficient for EEG data. The challenge is further exacerbated by the inherent difficulties associated with EEG data, which, even when ethically collected and cleaned, still suffer from limited data availability and significant variability across sessions and subjects.
Subject-wise analysis is essential in practice due to the inherently private and sensitive nature of EEG data. Models like ours, designed for thought-to-text translation, must be developed and deployed within privacy-preserving settings. Creating a personalized solution without access to extensive cross-subject EEG data can be challenging. To this end, our multi-stage approach demonstrates that pretraining on non-EEG data (such as images) and fine-tuning on small amounts of subject-specific EEG data opens new avenues for privacy-preserving personalized EEG-LLM model development.
@@ 6.4 Qualitative Inspection and Basic Error Analysis
For qualitative inspection, we compared the generated descriptions of images produced by both the GPT-4 and the Mistral-7B-v0.3 model (ALL and variants). Table 3 in the Appendix section presents notable examples from our assessment of the model’s generated descriptions. In many instances, the approach with input of EEG + predicted object label generates highly accurate descriptions, as illustrated in positive examples 2 and 4 for the piano and pumpkin, respectively. However, in cases of misgeneration, we encounter not only significantly inaccurate outputs and hallucinations, as seen in examples 6, 7, and 8, but also instances of genuine confusion. For example, the EEG signal for a flower is misidentified as a mushroom in example 5, leading to incorrect generation. This misidentification indicates potential areas for improvement in object classification.
However, a key advantage of the EEG-only approach is that even when the predicted object label is incorrect, the model can still produce reasonably coherent descriptions, as seen in anecdotal example 1. This highlights the robustness of the EEG embeddings in guiding the language model’s generation. Similar observations were made with other models based on LLaMa and Qwen architectures, further validating the consistency and reliability of our approach across various LLM frameworks. We acknowledge that the selected examples are purposefully chosen to illustrate different scenarios.
@@ 7 Conclusion and Future Work
In this paper, we introduced a novel approach to convert EEG signals into text, leveraging instruction-tuned LLMs fine-tuned with EEG data. Our method progresses through three stages: training an EEG encoder for feature extraction, fine-tuning LLMs on multimodal data, and further refining them with EEG embeddings for direct text generation from neural signals. Validation on a public EEG dataset demonstrates the efficacy of popular LLMs in "transforming" thoughts evoked by viewing images into text. We observed significant performance improvements compared to chance evaluation, and our methodology, incorporating all stages, performed well in both cross-subject and in-subject analyses, as validated through quantitative evaluation. The qualitative evaluation provided further insights into various scenarios involving EEG signals and object labels versus EEG signals alone as inputs for generating text. These evaluations not only reinforce the effectiveness of our methodology for efficient text generation but also underscore the potential of utilizing EEG data alone to achieve the desired results. However, instances of misidentification that result in incorrect outputs reveal opportunities for further improvements in text generation.
Our future work will focus on optimizing the model architecture, leveraging foundational pretrained EEG models on diverse dataset, improving EEG-text alignment through stage 2 training on large scale image datasets and diverse tasks such as optical character recognition, question answering and summarizing and exploring practical applications in healthcare and assistive technologies, marking a significant stride toward accessible "thoughts-to-text" systems.
@@
Extracting fine-grained information from EEG signals presents challenges due to high data-to-noise ratio and low spatial resolution. Despite these difficulties, EEG signals can still identify object categories, which can then be used with a generic prompt to aid in text generation. However, misclassification of similar-shaped objects (see the Appendix, Figure 3), such as mistaking a mushroom for a flower, underscores potential ambiguities in object recognition. In addition, there were instances where the encoder generated unrelated descriptions, such as identifying a coffee maker as a computer or an elephant as a panda. Object classification accuracy varies among subjects (see the Appendix, Table 4) due to individual differences in interpreting images, leading to diverse EEG signal variations and increased prediction variance. One approach to address this is by training personalized models for each subject and assessing their performance. Additionally, implementing methods that enhance the generalizability of predictions across subjects could be explored. Another challenge is data scarcity; Deep Learning models typically require substantial data for training. Hence, acquiring more high-quality, multi-channel EEG data under controlled experimental conditions is crucial to reduce noise. Despite these challenges, our quantitative and qualitative findings demonstrate promising results. We believe that additional training on larger datasets and rigorous controlled experiments will significantly improve performance.
@@
For this work, we utilized anonymized open-source EEG data, acknowledging the sensitivity of EEG data and the imperative of ethical compliance. All experimental data used in our research were anonymized to protect participant privacy and uphold ethical standards. Additionally, we employed OpenAI’s ChatGPT-4 system to enhance writing efficiency by generating LaTeX code, ensuring concise sentences, and aiding in error debugging.